Twist Company Stated operation in the year, 2019. The business assets are financed by both equity and debt. The values of debt and equity of the business had been represented by Debt: Equity ratio. The ratio of Debt: Equity of Twist Company is 1: 0.4. Additional Information Twist Company Ltd.’s dividend has a beta of 1.223. The risk-free rate is 10% and the market is expected to yield a return of 19%. The cost of debt is 10% and corporate tax is 25%. Required i. Compute the Cost of Debt after tax (2 marks) ii. Compute the value of Equity and Debt in Twist Company Limited (in terms of percentage) (3 marks) iii. Calculate the Cost of Equity Capital of Twist Company Limited (4 marks) iv. Calculate the Weighted Average Cost of Capital (WACC) of Twist Company Ltd. (5 marks) v. Determine the effect of an Increase in Tax Rate on the Weighted Average Cost of Capital (WACC) of Twist Company Ltd. (6 marks) vi. Explain why Debt is cheaper than Equity Capital (10 marks) (Total: 30 marks)
Answer #1
i] Cost of debt after tax = 10%*(1-25%) = 7.50%
ii] Value of debt = 1/1.4 = 71.43%
Value of equity = 0.4/1.4 = 28.57%
iii] Cost of equity per CAPM = 10%+1.223*(19%-10%) = 21.01%
iv] WACC = 7.50%*71.43%+21.01%*28.57% = 11.36%
v] Effect of WACC, when tax rate increases to 30%:
Cost of debt after tax = 10%*(1-30%) = 7.00%
WACC = 7.00%*71.43%+21.01%*28.57% = 11.00%
When the tax rate increases, the cost of debt and
thereby the WACC will decrease.
vi] Debt is cheaper due to two reasons:
*The suppliers of debt require lower returns, as debt
is a senior security in comparison with preferred stock
and common equity.
*The interest expense is tax deductible which, further
lowers the cost of debt.